Popper Corp. issues 8% bonds at 96 on January 1, 2019. The bonds have a face value of $10,000,000, pay coupons semiannually on June 30 and December 31, and mature in 10 years. Popper uses straight-line amortization for discounts and premiums. Accordingly, at December 31, 2021, the unamortized discount is $280,000. On October 1, 2022, Popper invokes a call option and extinguishes the bonds at 102 plus accrued interest. How much gain or loss should Popper recognize on the early extinguishment?

A. $637,500 LOSS

B. $450,000 LOSS

C. $200,000 LOSS

D. $400,000 LOSS
Answer #1

Answer : C) $450,000 Loss

Face Value of the Bond =$ 10,000,000

Issued At 96 , discount = 1-96 =4%

Discount on Bonds Payable = $10,000,000*4% = $400,000

bond Payable is for 10years, semi annula Bond Amortization = 10*2 = 20

Semi-annula Bond amortization = $400,000/20 = $20,000

Unamortized Bond discount on Decmeber 31, 2021 = $280,000

Bond Discount amortization from Jnauary 1st 2022 to oct 1,2022 = $20,000+ 20,000*3/6 = 30,000

Total Bond Carrying value on Oct 1st 2022 = $10,000,000- (280,000-30000) = $9,750,000

Bond retire at 102 = 4 10,000,000*1.02 = $10,200,000

Loss Ob bond Retire = $ 10,200,000- $ 9,750,000 = $450,000